Will someone please help me solve these problems? I belive I have them correct, but will you check my answers Please.
1)Factor: 36x^2-z^2 = (6x+z)(6x-v)
2) Factor the quadatric expressions: w^2+6w+9 +(w+3) (w=3)
3) factor completely: 4x^3-10x^2-24x=(x-4)(2+3)
4) Factor: 2x^2+3xy-20y^2

Answer by mananth(16949)

1)Factor: 36x^2-z^2 = (6x+z)(6x-z)
''
2) w^2+6w+9 =(w+3)(w+3)
..
3) factor completely: 4x^3-10x^2-24x
=2x(2x^2-5x-12)
=2x(2x^2-8x+3x-12)
=2x((2x(x-4)+3(x-4))
=2x(2x-4)(2x+3)
..
4) 2x^2+3xy-20y^2
=2x^2+8xy-5xy-20y^2
=2x(x+4y)-5y(x+4y)
=(x+4y)(2x-5y)